OpenStudy (ashy98):

A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = distance over time; 1 mile = 1609 meters. A.) 10.4 meters per second B.) 23.3 meters per second C.) 52.2 meters per second D.) 93.9 meters per second

OpenStudy (ashy98):

okay so it would be A

OpenStudy (anonymous):

you divide 70 / 1609 then you get the answer :D thats all. its a simple division problem :D multiple choices are my specialty :D

OpenStudy (dangerousjesse):

You use \(70\times 1609=x, x\div 3=y, y\div 120= the~~answer~~\)

OpenStudy (anonymous):

wait what? :/ hold on didnt see that

OpenStudy (dangerousjesse):

@FweaBoyLive If the bird flies 1600 meters for every mile, that's 1609*70 :)

OpenStudy (wolf1728):

70 miles / 3 hours = 23.33333 miles / hour 23.333333 miles x 1609 meters / hour = 37543.33 meters / hour = 37,543.33 / 3,600 = 10.429 meters / second

OpenStudy (anonymous):

oh i had it right nvm

OpenStudy (ashy98):

thank you guys soooo much!

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

you can divide it top its the short way :D

OpenStudy (anonymous):

too*